How would I solve
If a number is increased by 5 and then trebled, the result is six more than two thirds of the number. Find the number.
I have
(x+5)^3=2|3x+6

Answer by edjones(7569)

Let the number be x.
3(x+5) = 2x/3 + 6
9(x+5)=2x+18 Multiply each side by 3
9x+45=2x+18
7x=-27
x=-27/7
.
Ed